multiply both side of the equation by a common denominator to eliminate the
fractions. Then solve the system.

x/4+3y=31
2x-y/5=8
I have been struggling to try to solve this for over 4 days now and have had no luck. We are given the final answer which is (5,10) but I can not figurge out how they got to that. I need to show the work and solve the problem. Can someone please help me.

Answer by mananth(16949)

while writing the problem it is advisable to put brackets to get clarity.
x/4+3y=31 ...............1
2x-y/5=8 ................2
...
multiply by 4
4*x/4 +4*3y= 31*4
x+12y=124...........3
..
2x- y/5 =8
multiply by 5
5*2x-5*y/5= 5*8
10x-y= 40.............4
..
..
multiply equation 3 by -10 and to equation 4
-10x-120y+10x-y=-1240+40
-121y=-1200
y= 1200/121
plug the value of y in the above equation
x=604/121